OpenStudy (anonymous):

Let u = <-5, -9>, v = <6, 8>. Find -8u - 2v.

OpenStudy (anonymous):

<52, 88> <-8, 2> <112, -28> <28, 56>

OpenStudy (anonymous):

\[u \in (-5,-9)\]\[v \in (6,8)\] for fining the minimum of -8u - 2v we choose the smallest number of intervals. -8*-5 - 2*6 =40-12=28 for fining the maximum of -8u - 2v we choose the largest number of intervals. -8*-9 - 2*8 =72-16=56 so (28, 56) is the answer